# Designing Phase Sensitive Probes of Monopole Superconducting Order

### v1

@@ IV.1.2 Method 2: Numerical results from tight-binding model
We further show the nonzero Josephson coupling in the same Josephson junction between an $s$-wave superconductor and $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor, as shown in Fig. 2 (a), using a tight-binding model on a cubic lattice (see Appendix D and E for details). As $k_{x}$ and $k_{y}$ are assumed to be conserved across the junction, we take periodic boundary conditions in the $x$ and $y$ directions. We consider $2N_{z}$ sites along $z$ direction and take open boundary conditions at the two ends at $n_{z}=1$ and $2N_{z}$. The junction barrier is located between $n_{z}=N_{z}$ and $n_{z}=N_{z}+1$, in which only spin-independent nearest neighbor hopping within the barrier is considered. At given conserved momenta $k_{x}$ and $k_{y}$, we model the barrier along $z$ direction as
$H_{\mathrm{link};k_{x},k_{y}}=t_{0}\sum_{\sigma=\uparrow,\downarrow}(c^{\dagger}_{k_{x},k_{y};\sigma,N_{z}+1}c_{k_{x},k_{y};\sigma,N_{z}}+\mathrm{h.c.})$ (33)
@@
in which $t_{0}$ is the spin-independent hopping amplitude, and $c_{k_{x},k_{y};\sigma,n_{z}}$ annihilates a particle at site $n_{z}$ with momentum $k_{x}$ and $k_{y}$ and spin $\sigma$. If we increase the thickness of the barrier to a few sites, there is no qualitative change in Josephson current.
@@
Figure 3: Numerical results for a tight-binding model of the Josephson junction along $z$ direction between an $s$-wave superconductor and $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor. (a) Maximal cross section of bulk Fermi surface $\mathrm{FS}_{+}$ of the monopole superconductor at $k_{z}=K_{0,z}$ (solid) and that of the Fermi surface $\mathrm{FS}_{0}$ of the $s$-wave superconductor at $k_{z}=0$ (dashed). Here, $\mu_{L}/t_{0}=3.88$ and $\mu_{R}/t_{0}=0.05$. (b) The Josephson energy phase relation contributed by different momenta in (a), $k_{\parallel,n}=(n-1)\tilde{k}_{F}/4$, $n=1,\cdots,5$, and $\tilde{k}_{F}=0.2$. Here, the relative magnitude is given by $\Delta^{(n)}_{0}=\Delta_{0}k_{\parallel,n}/\tilde{k}_{F}$, with $\Delta_{0}/t=0.04$. Lowest energy states for a given radial cut correspond to the cuts shown in (a) and are shown in color, while the gray bands represent higher-energy contributions. (c) The total Josephson current phase relation (red, solid) compared to that of a $z$ direction junction between a $p_{x}-ip_{y}$ superconductor and an $s$-wave SC (black, dashed). (d) The Fraunhofer pattern of a corner junction (inset) in the $xy$-plane of a monopole superconductor (red, solid), compared to that of a $p_{x}+ip_{y}$ superconductor (black, dashed). Above, $I_{0}={e\Delta_{0}}/{\hbar}$, and the system size is $3002N_{z}=$ 300.
@@
In Fig. 3, we present our numerical results of the Josephson current for the junction between an $s$-wave superconductor and a $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor. Fermi surfaces of the bulk tight-binding Hamiltonians at the two sides of the junction are shown in Fig. 3 (a). Although the lattice has four-fold rotation symmetry with respect to $z$ axis, the Fermi surfaces are nearly spherical in the $k_{x}k_{y}$ plane. In Fig. 3 (b), contributions to the Josephson energy phase relation are calculated for representative momenta ${\mathbf{k}}_{\parallel}$ at different radii $k_{\parallel}=\sqrt{{{k}_{x}^{2}+{k}_{y}^{2}}}$. At a given value of $k_{\parallel}$, different azimuthal angles, $\varphi_{\tilde{k}}$, contribute nearly the same energy phase relation. Hence, the contributions from states at two representative azimuthal angles, $\varphi_{\tilde{k}}=0$ and $\pi/4$, shown in the inset of Fig. 3 (a), overlap with each other in Fig. 3 (b). At $k_{\parallel}=k_{\parallel,1}=0$, the states close to north and south poles of the Fermi surfaces contribute to the Josephson coupling. However, the magnitude of the pairing gap of the monopole superconductor, $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$, vanishes at the poles of the Weyl Fermi surfaces. As such, states at $k_{\parallel,1}$ correspond to vanishingly small Josephson energy and therefore flat energy phase relation. As $k_{\parallel}$ increases to nonzero values of $k_{\parallel,i}$ with $i\geq 2$, the corresponding contribution to Josephson energy shows significant $2\pi$ periodicity. Furthermore, the overall magnitude of Josephson energy increases as the magnitude of pairing order $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ at $k_{\parallel,i}$ increases until reaching its maximum at $k_{\parallel,5}\approx\tilde{k}_{F}$.
@@
In addition to the net $2\pi$-periodic Josephson energy phase relation, there are contributions from second order Josephson tunneling, where multiple scatterings from the junction interface lead to $\pi$-periodic contributions to the energy phase relation. Such higher order contributions are most significant for states near the gap nodes of the $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole pairing order, for example, at $k_{\parallel}=k_{\parallel,2}$ in Fig. 3 (b). We provide further detailed explanation on this in Appendix B.
@@
We calculate the total Josephson current at zero temperature, $\left.I_{J}(\phi)\right|_{T=0}=(-{2e}/{\hbar})\sum_{\mathbf{k}_{\parallel}}{\partial E_{J}(\mathbf{k}_{\parallel},\phi)}/{\partial\phi}.$ Here, $E_{J}(\mathbf{k}_{\parallel},\phi)$ is the Josephson energy phase relation for occupied states with transverse momentum $\mathbf{k}_{\parallel}$. In Fig. 3 (c), we show our numerical result of Josephson current with the summation over transverse momentum $k_{\parallel}\leq\tilde{k}_{F}$ at representative values, $k_{\parallel,i=1,\cdots,5}$, and taking account of the measure in polar coordinates. For $k_{\parallel}>\tilde{k}_{F}+{\Delta_{R,0}}/{v_{F}}$, there are no in-gap Andreev states contributing to the Josephson current as discussed in Sec. IV.1.1. Moreover, comparing with Fig. 2 (c), the total Josephson current phase relation in Fig. 3 (c) is modulated by $\pi$-periodic contributions from higher order tunneling processes. However, the total Josephson current, obtained by superimposing contributions from different $k_{\parallel}$, is nonetheless $2\pi$-periodic, as described below.
@@
The first order $2\pi$-periodic Josephson current between an $s$-wave superconductor and $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor is nonvanishing in this tight-binding model for the same symmetry reasons discussed for the continuum model in Sec. IV.1.1. The tight-binding model preserves the two parity-related Fermi surfaces which have opposite Chern numbers. Although the Fermi surfaces are not perfectly spherical, the topology of the effective pairing remains invariant and is still classified by the pair monopole charge, $q_{p}=-1$. Additionally, the global angular momentum of the pairing order is constrained to the $l_{z}=0$ channel, as both the kinetic and the proximitized pairing parts of the tight-binding Hamiltonian are invariant under lattice rotations $R_{z}$. Therefore, the effective pairing order likewise transforms according to conserved $l_{z,\mathrm{glob}}=0$. However, compared to the effective pairing derived in the continuum model in Eq. (24), there are small but nonzero contributions to the pairing order from monopole harmonics in the $l=2$ partial wave channel as a result of the lattice symmetry. Nonetheless, these contributions from higher partial wave channels do not significantly affect the nonzero first order Josephson coupling. Hence, because the tight binding model preserves the symmetry and topology of the system, the essential features of the Josephson current, for example its $2\pi$-periodicity, agree with the previous result based on continuum model.
@@
We additionally examine a corner junction between a $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor and $s$-wave superconductor. The corner junction is symmetrically aligned with two faces along the $x$ and $y$ directions, as shown schematically in the inset in Fig. 3 (d). Because this geometry is sensitive to pairing phase difference along the two orthogonal crystal axes [20], it can be used as an additional probe of global angular momentum $l_{z,\mathrm{glob}}$. Based on the same tight-binding model of $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor, we study the resulting critical current in the presence of an external magnetic field along the $z$ direction. As the external flux $\Phi_{B}$ through the barrier of the corner junction varies, we obtain the Fraunhofer pattern of the corner junction, as shown in Fig. 3 (e). The Fraunhofer pattern is qualitatively the same as that of an SIS junction between two $s$-wave superconductors, with nodes at integer flux quantum, $\Phi_{0}=h/2e$. This is in contrast to, for example, a corner junction between an $s$-wave and chiral $p$-wave superconductor, which gives rise to an interference pattern that is asymmetric with respect to zero flux.
@@
In alignment with the findings presented in the preceding section, in which the Josephson current was analyzed through linear response theory, the results from the tight-binding model further reinforce that, globally, the $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconducting pairing order transforms according to $l_{z,\mathrm{glob}}=0$. Consequently, the monopole superconductor has nonzero Josephson coupling with an $s$-wave superconductor in a junction along the $z$ axis, leading to an overall nonvanishing $2\pi$-periodic first order Josephson current. We next discuss similar junction designs to probe monopole superconductors with nonzero global angular momentum $z$ component, or when the pairing order is in a different topological sector.
@@ IV.2 Probing monopole superconducting orders with nonzero global angular momenta $l_{z,\mathrm{glob}}$
We now demonstrate that the $z$ direction Josephson junction design introduced in Sec. III.2 can be generalized to probe monopole pairing orders with nonzero $z$-component global angular momentum. As an example, we consider the monopole pairing order in the same topological sector as that in Sec. IV.1 given by the pair monopole charge $q_{p}=-1$, but in a different angular momentum channel with $l=l_{z}=2$. Following the symmetry argument, the first order Josephson coupling is allowed only when coupled to another superconductor with pairing order in the $l_{z}=2$ channel, for example, a $d_{x^{2}-y^{2}}+id_{xy}$ wave superconductor, in an SIS junction with geometry preserving the rotational symmetry about $z$ axis. To demonstrate the symmetry principle, here, we focus on deriving the form factor $\mathfrak{F}_{\mathrm{sing}-\mathrm{MSC}}^{(q_{p}=-1)}$ in this system and show that it leads to nonzero first order Josephson coupling. In Appendix F, we also show nonvanishing first order Josephson current obtained numerically from a tight-binding model of this system.
@@
where, in consistency with prior notation, $\lambda=\pm 1$ corresponds to a gauge in which the Dirac string is located at $\theta_{\tilde{k}}=\pi(1+\lambda)/2$. As the Chern numbers of the Fermi surfaces that participate in the inter-Fermi surface pairing are unchanged, the effective pairing order remains in the same topological sector characterized by the pair monopole charge, $q_{p}=-1$. However, the pairing order now transforms according to its global angular momentum, $l_{z}=l_{z,\mathrm{glob}}=2$. The nonzero global angular momentum can be attributed to the inter-Fermi surface $(p_{x}+ip_{y}){|\uparrow\uparrow\rangle}$ pairing, which transforms according to its total angular momentum $j_{z}=2$ in the presence of spin-orbit coupling. When coupled to the spin-texture of the Weyl semimetal Fermi surface, this induces the global $l_{z}=2$ winding of the monopole superconducting order in momentum space.
@@
with $\Delta_{R,0}$ being the amplitude of the inter-Fermi surface pairing. Because both the $d_{x^{2}-y^{2}}+id_{xy}$ superconductor and $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=2)}$ monopole superconductor transform under rotation $R_{z}$ according to their global angular momentum $l_{z}=2$, the form factor itself is invariant under $R_{z}$. Hence, per the same symmetry arguments in Sec. IV.1.1, this leads to nonvanishing first order Josephson coupling.

### v2

@@ IV.1.2 Method 2: numerical results from tight-binding model
We further show the nonzero Josephson coupling in the same Josephson junction between an $s$-wave superconductor and $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor, as shown in Fig. 2(a), using a tight-binding model on a cubic lattice (see Appendix D and E for details). As $k_{x}$ and $k_{y}$ are assumed to be conserved across the junction, we take periodic boundary conditions in the $x$ and $y$ directions. We consider $2N_{z}$ sites along $z$ direction and take open boundary conditions at the two ends at $n_{z}=1$ and $2N_{z}$. The junction barrier is located between $n_{z}=N_{z}$ and $n_{z}=N_{z}+1$, where only spin-independent nearest neighbor hopping within the barrier is considered. At given conserved momenta $k_{x}$ and $k_{y}$, we model the barrier along $z$ direction as
$H_{\mathrm{link}}(k_{x},k_{y})=t_{0}\sum_{\sigma=\uparrow,\downarrow}(c^{\dagger}_{k_{x},k_{y};\sigma,N_{z}+1}c_{k_{x},k_{y};\sigma,N_{z}}+\mathrm{h.c.})$ (33)
@@
where $t_{0}$ is the spin-independent hopping amplitude, and $c_{k_{x},k_{y};\sigma,n_{z}}$ annihilates a particle at site $n_{z}$ with momentum $k_{x}$ and $k_{y}$ and spin $\sigma$. If we increase the thickness of the barrier to a few sites, there is no qualitative change in Josephson current.
@@
Figure 3: Numerical results for a tight-binding model of the Josephson junction along $z$ direction between an $s$-wave superconductor and $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor. (a) Maximal cross section of bulk Fermi surface $\mathrm{FS}_{+}$ of the monopole superconductor at $k_{z}=K_{0,z}$ (solid) and that of the Fermi surface $\mathrm{FS}_{0}$ of the $s$-wave superconductor at $k_{z}=0$ (hatched). Here, $\mu_{L}/t_{0}=3.88$ and $\mu_{R}/t_{0}=0.05$. (b) The Josephson energy phase relation contributed by different momenta in (a), $k_{\parallel,n}=(n-1)\tilde{k}_{F}/4$, $n=1,\cdots,5$, and $\tilde{k}_{F}=0.2$. Here, the relative magnitude is given by $\Delta^{(n)}_{0}=\Delta_{0}k_{\parallel,n}/\tilde{k}_{F}$, with $\Delta_{0}/|t_{0}|=0.04$. Lowest energy states for a given radial cut correspond to the cuts shown in (a) and are shown in color, while the gray bands represent higher-energy contributions. (c) The total Josephson current phase relation (red, solid) compared to that of a $z$ direction junction between a $p_{x}-ip_{y}$ superconductor and an $s$-wave SC (black, dashed). (d) The Fraunhofer pattern of a corner junction (inset) in the $xy$-plane of a monopole superconductor (red, solid), compared to that of a $p_{x}+ip_{y}$ superconductor (black, dashed). Here, $I_{0}={e\Delta_{0}}/{\hbar}$, and the system size is $3002N_{z}=$ 300.
@@
In Fig. 3, we present our numerical results of the Josephson current for the junction between an $s$-wave superconductor and a $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor. Since the junction breaks translational symmetry along $z$ direction, longitudinal momentum $k_{z}$ is not conserved. This effectively allows Cooper pairs with different $k_{z}$ components on either side of the junction to tunnel across the interface. Consequently, even with mismatched $3$D Fermi surfaces, a nonzero Josephson current can still occur as long as their $2$D projections in the transverse momentum plane overlap sufficiently, as illustrated in Fig. 3(a). Although the lattice has fourfold rotation symmetry with respect to the $z$ axis, the projected Fermi surfaces are nearly spherical in the $k_{x}k_{y}$ plane. In Fig. 3(b), contributions to the Josephson energy phase relation are calculated for representative momenta ${\mathbf{k}}_{\parallel}$ at different radii $k_{\parallel}=\sqrt{{{k}_{x}^{2}+{k}_{y}^{2}}}$. At a given value of $k_{\parallel}$, different azimuthal angles, $\varphi_{\tilde{k}}$, contribute nearly the same energy phase relation. Hence, the contributions from states at two representative azimuthal angles, $\varphi_{\tilde{k}}=0$ and $\pi/4$, shown in the inset of Fig. 3(a), overlap with each other in Fig. 3(b). At $k_{\parallel}=k_{\parallel,1}=0$, the states close to north and south poles of the Fermi surfaces contribute to the Josephson coupling. However, the magnitude of the pairing gap of the monopole superconductor, $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$, vanishes at the poles of the Weyl Fermi surfaces. As such, states at $k_{\parallel,1}$ correspond to vanishingly small Josephson energy and therefore flat energy phase relation. As $k_{\parallel}$ increases to nonzero values of $k_{\parallel,i}$ with $i\geq 2$, the corresponding contribution to Josephson energy shows significant $2\pi$ periodicity. Furthermore, the overall magnitude of Josephson energy increases as the magnitude of pairing order $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ at $k_{\parallel,i}$ increases until reaching its maximum at $k_{\parallel,5}\approx\tilde{k}_{F}$.
@@
In addition to the net $2\pi$-periodic Josephson energy phase relation, there are contributions from second-order Josephson tunneling, where multiple scatterings from the junction interface lead to $\pi$-periodic contributions to the energy phase relation. Such higher-order contributions are most significant for states near the gap nodes of the $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole pairing order, for example, at $k_{\parallel}=k_{\parallel,2}$ in Fig. 3(b). We provide further detailed explanation on this in Appendix B.
@@
We calculate the total Josephson current at zero temperature, $\left.I_{J}(\phi)\right|_{T=0}=(-{2e}/{\hbar})\sum_{\mathbf{k}_{\parallel}}{\partial E_{J}(\mathbf{k}_{\parallel},\phi)}/{\partial\phi}.$ Here, $E_{J}(\mathbf{k}_{\parallel},\phi)$ is the Josephson energy phase relation for occupied states with transverse momentum $\mathbf{k}_{\parallel}$. In Fig. 3(c), we show our numerical result of Josephson current with the summation over transverse momentum $k_{\parallel}\leq\tilde{k}_{F}$ at representative values, $k_{\parallel,i=1,\cdots,5}$, and taking account of the measure in polar coordinates. For $k_{\parallel}>\tilde{k}_{F}+{\Delta_{R,0}}/{v_{F}}$, there are no in-gap Andreev states contributing to the Josephson current as discussed in Sec. IV.1.1. Moreover, comparing with Fig. 2(c), the total Josephson current phase relation in Fig. 3(c) is modulated by $\pi$-periodic contributions from higher-order tunneling processes. However, the total Josephson current, obtained by superimposing contributions from different $k_{\parallel}$, is nonetheless $2\pi$-periodic.
The Josephson current can be further enhanced when the Weyl points of the bulk monopole superconductor are close to the Fermi surface of the bulk s-wave superconductor. As an additional example, we keep the overlap area of the $2$D projections of the Fermi surfaces in the transverse momentum plane identical to that in Fig. 3(a). However, we reduced the $s$-wave superconductor’s Fermi wavevector from $k_{F}=1.63$ ($\mu_{L}/t_{0}=3.88$) to $k_{F}=1.05$ ($\mu_{L}/t_{0}=5$), ensuring that the bulk Fermi surfaces at two sides of the junction also overlap in $3$D momentum space. We observed $28$% enhancement in the critical Josephson current. Nonetheless, the qualitative features remained unchanged, including the nonvanishing first-order Josephson current as well as the periodicity of the Josephson current phase relations.
@@
the first-order $2\pi$-periodic Josephson current between an $s$-wave superconductor and $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor is nonvanishing in this tight-binding model for the same symmetry reasons discussed for the continuum model in Sec. IV.1.1. The tight-binding model preserves the two parity-related Fermi surfaces which have opposite Chern numbers. Although the Fermi surfaces are not perfectly spherical, the topology of the effective pairing remains invariant and is still classified by the pair monopole charge, $q_{p}=-1$. Additionally, the global angular momentum of the pairing order is constrained to the $l_{z}=0$ channel, as both the kinetic and the proximitized pairing parts of the tight-binding Hamiltonian are invariant under lattice rotations $R_{z}$. Therefore, the effective pairing order likewise transforms according to conserved $l_{z,\mathrm{glob}}=0$. However, compared to the effective pairing derived in the continuum model in Eq. (24), there are small but nonzero contributions to the pairing order from monopole harmonics in the $l=2$ partial wave channel as a result of the lattice symmetry. Nonetheless, these contributions from higher partial wave channels do not significantly affect the nonzero first-order Josephson coupling. Hence, because the tight-binding model preserves the symmetry and topology of the system, the essential features of the Josephson current, for example its $2\pi$-periodicity, agree with the previous result based on continuum model. Furthermore, these features persist in the long Josephson junction limit, as discussed in Sec. VI.
@@
We additionally examine a corner junction between a $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor and $s$-wave superconductor. The corner junction is symmetrically aligned with two faces along the $x$ and $y$ directions, as shown schematically in the inset in Fig. 3(d). Because this geometry is sensitive to pairing phase difference along the two orthogonal crystal axes Van Harlingen (1995), it can be used as an additional probe of global angular momentum $l_{z,\mathrm{glob}}$. Based on the same tight-binding model of $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconductor, we study the resulting critical current in the presence of an external magnetic field along the $z$ direction. As the external flux $\Phi_{B}$ through the barrier of the corner junction varies, we obtain the Fraunhofer pattern of the corner junction, as shown in Fig. 3(d). The Fraunhofer pattern is qualitatively the same as that of an SIS junction between two $s$-wave superconductors, with nodes at integer flux quantum, $\Phi_{0}=h/2e$. This is in contrast to, for example, a corner junction between an $s$-wave and chiral $p$-wave superconductor, which gives rise to an interference pattern that is asymmetric with respect to zero flux.
@@
In alignment with the findings presented in the preceding section, where the Josephson current was analyzed through linear response theory, the results from the tight-binding model further reinforce that, globally, the $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=0)}$ monopole superconducting pairing order transforms according to $l_{z,\mathrm{glob}}=0$. Consequently, the monopole superconductor has nonzero Josephson coupling with an $s$-wave superconductor in a junction along the $z$ axis, leading to an overall nonvanishing $2\pi$-periodic first-order Josephson current. We next discuss similar junction designs to probe monopole superconductors with nonzero global angular momentum $z$ component, or when the pairing order is in a different topological sector.
@@ IV.2 Probing monopole superconducting orders with nonzero global angular momenta $l_{z,\mathrm{glob}}$
We now demonstrate that the $z$ direction Josephson junction design introduced in Sec. III.2 can be generalized to probe monopole pairing orders with nonzero $z$-component global angular momentum. As an example, we consider the monopole pairing order in the same topological sector as that in Sec. IV.1 given by the pair monopole charge $q_{p}=-1$, but in a different angular momentum channel with $l=l_{z}=2$. Following the symmetry argument, the first-order Josephson coupling is allowed only when coupled to another superconductor with pairing order in the $l_{z}=2$ channel, for example, a $d_{x^{2}-y^{2}}+id_{xy}$ wave superconductor, in an SIS junction with geometry preserving the rotational symmetry about $z$ axis. To demonstrate the symmetry principle, here, we focus on deriving the form factor $\mathfrak{F}_{\mathrm{sing}-\mathrm{MSC}}^{(q_{p}=-1)}$ in this system and show that it leads to nonzero first-order Josephson coupling. In Appendix F, we also show nonvanishing first-order Josephson current obtained numerically from a tight-binding model of this system.
@@
where, in consistency with prior notation, $\lambda=\pm 1$ corresponds to a gauge where the Dirac string is located at $\theta_{\tilde{k}}=\pi(1+\lambda)/2$. As the Chern numbers of the Fermi surfaces that participate in the inter-Fermi surface pairing are unchanged, the effective pairing order remains in the same topological sector characterized by the pair monopole charge, $q_{p}=-1$. However, the pairing order now transforms according to its global angular momentum, $l_{z}=l_{z,\mathrm{glob}}=2$. The nonzero global angular momentum can be attributed to the inter-Fermi surface $(p_{x}+ip_{y}){|\uparrow\uparrow\rangle}$ pairing, which transforms according to its total angular momentum $j_{z}=2$ in the presence of spin-orbit coupling. When coupled to the spin-texture of the Weyl semimetal Fermi surface, this induces the global $l_{z}=2$ winding of the monopole superconducting order in momentum space.
@@
with $\Delta_{R,0}$ being the amplitude of the inter-Fermi surface pairing. Because both the $d_{x^{2}-y^{2}}+id_{xy}$ superconductor and $\Delta_{\mathrm{MSC}}^{(q_{p}=-1,l_{z}=2)}$ monopole superconductor transform under rotation $R_{z}$ according to their global angular momentum $l_{z}=2$, the form factor itself is invariant under $R_{z}$. Hence, per the same symmetry arguments in Sec. IV.1.1, this leads to nonvanishing first-order Josephson coupling.